OpenStudy (anonymous):

one number plus five times its reciprocal is equal to six. find the number or numbers.

OpenStudy (anonymous):

x(1/x) = 6

OpenStudy (anonymous):

Does that make sense? The reciprocal of a number is anything over it Now you have to simplify x(1/x) = x/x Oh wait I'm lost now x/x = 6 x would have to be equal to two numbers. Is that possible? What did I do wrong

OpenStudy (lieutenantgeneral):

Suppose the number is x. Then x + 5/x = 6 Multiplying throughout by x, x2 + 5 = 6x or x2 - 6x + 5 = 0 which can be factorised into (x-1)*(x-5)=0 So x = 1 (and 5 times its reciprocal is 5) or x = 5 (and 5 times its reciprocal is 1).

OpenStudy (anonymous):

Wait no it's not x + 5(1/x)